OpenStudy (howard-wolowitz):

Identify the middle line for the function:

OpenStudy (jdoe0001):

ahemmm I'd say, graph it :)

OpenStudy (howard-wolowitz):

i did and nothing came up

OpenStudy (jdoe0001):

well... is the secant graph really with a few shifts here and there for the midline, the relevant shift is the vertical shift though \(\textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards \\ \quad \\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ % template start f(t) = {\color{purple}{ 5}}sec ( {\color{blue}{ 3}}t + {\color{red}{ \pi }} ) + {\color{green}{ 1}}\\ % template ends \qquad\qquad\quad\ \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to\ the\ left \\ \quad \\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to\ the\ right\end{array} \end{array}\)

OpenStudy (jdoe0001):

hmmmm shold have been \(\pi\) rather than theta there anyhow, that's just a horizontal shift anyway, won't afffect the midline

OpenStudy (howard-wolowitz):

well how do u get a middle point

OpenStudy (howard-wolowitz):

so its A